OpenStudy (anonymous):

3r + - 40r = 111 please show all work

OpenStudy (zzr0ck3r):

factor the r out

OpenStudy (anonymous):

3r + - 40r = 111 -37r=111 -37r/-37=111/-37 r=-3

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

:)